OpenStudy (anonymous):

(sinA+sinB)^2+(cosA+cosB)^2-2 write as a single trig function

OpenStudy (unklerhaukus):

expand those squares with (x+y)^2 = x^2+2xy+y^2

OpenStudy (unklerhaukus):

then use sin^2 θ + cos^2 θ =1

OpenStudy (unklerhaukus):

most of the terms will cancel out

OpenStudy (unklerhaukus):

finally, use cos u cos v + sin u sin v = sin (u + v)

OpenStudy (whpalmer4):

@UnkleRhaukus Uh, that last identity doesn't look right to me. Try it with \( u = v = \pi/6\) I think \[\cos u \cos v + \sin u \sin v = \cos (u-v)\]is what you want there...

OpenStudy (unklerhaukus):

your right @whpalmer4

OpenStudy (anonymous):

im not sure what my deal is still not getting it the whole thing about trig id's

OpenStudy (whpalmer4):

Can you show us what you've done so far?

OpenStudy (anonymous):

2-(sinA-cosB)^2 -(cosA+sinB)^2 i expand the sq of both then i combine like terms and now im left with:2-sin^2A+cos^2B - cos^2A+2cosAsinB+Sin^2B

OpenStudy (whpalmer4):

It looks like you're starting in the wrong place: \[(\sin A+ \sin B)^2+(\cos A+\cos B)^2-2\]is what the problem statement reads.

OpenStudy (anonymous):

thats the problem out my book

OpenStudy (whpalmer4):

Well, that's not what you put up top!

OpenStudy (whpalmer4):

But no matter. You work this problem in much the same way. Expand the squares of both and you should get \[-2 \cos (x) \sin (y)+2 \sin (x) \cos (y)-\sin ^2(x)-\cos ^2(x)-\sin ^2(y)-\cos ^2(y)+2\] Use \(\sin^2(x) + \cos^2(x)= 1\) to simplify that (also do it for \(y\). When you are done, you'll be left with the \(\sin * \cos\) product terms, and you should be able to find them on your trig identity sheet.

OpenStudy (whpalmer4):

where did you get that from?

OpenStudy (anonymous):

oops the problem up top is a diff. i dont see how -sin^2A-cos^2A=-1 i kno sin^2a+cos^a=1

OpenStudy (whpalmer4):

\[-\sin^2x -\cos^2x = -1\]Factor out a -1 \[-1(\sin^2x+\cos^2x) = -1\]\[-1(1) = -1\]

OpenStudy (anonymous):